OpenStudy (anonymous):

simplify the expression: 2sin^2x-1/sinx-cosx

OpenStudy (anonymous):

i always get confused with the 2 in front

OpenStudy (anonymous):

Is this the right way to rewrite your question? \[2\sin^2\theta-1/\sin(\theta)-\cos(\theta)\] Where the "1" divides both sine and cosine?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i'm not sure what to do

OpenStudy (anonymous):

Same here. Do you have any useful/relevant identities?

OpenStudy (zepp):

Hint: 1 = sin^2 (x) + cos^2 (x)

OpenStudy (anonymous):

i think its one of the pythagorean identities

OpenStudy (anonymous):

so what does the 2 in front do?

OpenStudy (zepp):

Something really interesting :) \[2\sin^2 x - 1\] If 1 = sin^2 (x) + cos^2 (x) \[2\sin^2(x) - (\sin^2x + \cos^2x)\] Then RELEASE the parentheses!!1! :D

OpenStudy (lgbasallote):

hint: it's gonna be like 2x - x where x = sin^2 x

OpenStudy (anonymous):

still confused...

OpenStudy (lgbasallote):

\[\large \frac{2sin^{2}x - sin^{2}x - cos^{2}x}{sinx - cosx}\] like i said 2sin^2 x - sin^ x is like 2x - x where x = sin^2 x when given 2x - x..what do you do??

OpenStudy (anonymous):

idk >.<' i'm still lost

OpenStudy (lgbasallote):

hmm maybe i shouldnt have used x...lemme try again..you have 2sin^2 x - sin^2 x right? let a = sin^2 x so... 2a - a = a right? and a = sin^2 x so...2 sin^2 x - sin^2 x = sin^2 x getting it now? :D

OpenStudy (anonymous):

not really... _ _||| i'm not really good at this

OpenStudy (lgbasallote):

hmmm..then i'll do this the traditional way 2sin^2 x - sin^2 x... imagine sin^2 x is an apple 2 apples - 1 apple...what do you have?

OpenStudy (anonymous):

1 apple

OpenStudy (lgbasallote):

and apple = sin^2 x so 1 apple = 1 (sin^2 x)..anything multiplied to 1 is the number itself so it's...sin^2 x do you get it now?

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

 is this what it looks like factored out?